The water level at an ocean inlet has a depth, d in meters, that varies with the time, t, in hours after midnight, according to the equation d = sin(2π(t-4/12.4)) + 6. What is the water depth at 5:30 ? To the nearest hundredth of a metre.
Answer by ikleyn(52898)

.
The water level at an ocean inlet has a depth, d in meters, that varies with the time, t,
in hours after midnight, according to the equation d = sin(2π(t-4/12.4)) + 6.
What is the water depth at 5:30 ? To the nearest hundredth of a metre.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

I will assime that 5:30 in this problem is 5:30 am;  so t= 5.5 hours  in the formula.


Then to get the answer, you should substitute t= 5.5 into the given formula


    d = sin(2pi*((5.5-4)/12.4))) + 6 = 6.69 meters  (approximately).    ANSWER

This problem is a matter of substituting a value into the formula.


Notice that in your formula in the post, a pair of parentheses is missed,
so I added it based on my common sense and the rules of Math grammar.